# On sharpened singular Adams’ type inequalities and applications to $(p,\frac{n}{2})$-biharmonic equations

### v2

@@ Abstract.
The purpose of this article is to establish a sharp version of Adams’ type inequality in a suitable higher order function space with singular weight in $\mathbb{R}^{n}$. In addition, we also provide the proof of a sharp singular concentration-compactness principle due to Lions’ as an improvement of this singular Adams’ inequality. It is well known that such types of inequalities can be classified in two different ways, for instance, critical sharp singular Adams’ type inequality (involving Sobolev full norm as constraint) and subcritical sharp singular Adams’ type inequality (involving Sobolev partial norm as constraint). Further, we shall establish that critical and subcritical sharp singular Adams’ type inequalities are surprisingly equivalent. To be more precisely, we also discuss about the asymptotic behavior of the lower and upper bounds of subcritical sharp singular Adams’ type inequality and establish a relation between the suprema of such types of critical and subcritical sharp inequalities. Despite this, we shall demonstrate a new compact embedding, which plays a crucial role in our arguments. Moreover, as an application of these results, by employing the mountain pass theorem, we study the existence of nontrivial solutions to a class of nonhomogeneous quasilinear elliptic equations involving $(p,\frac{n}{2})$-biharmonic operator with singular exponential growth as follows
$\Delta^{2}_{p}u+\Delta^{2}_{\frac{n}{2}}u=\frac{g(x,u)}{|x|^{\gamma}}\quad\text{in}\quad\mathbb{R}^{n},$
with $n\geq 4$, $1<p<\frac{n}{2}$, $\gamma\in(0,n)$ and the nonlinear term $g:\mathbb{R}^{n}\times\mathbb{R}\to\mathbb{R}$ is a Carathéodory function, which behaves like $\exp{(\alpha|s|^{\frac{n}{n-2}})}$ as $|s|\to~{}+\infty$ for some $\alpha>0$.
@@
$(p,\frac{n}{2})$-biharmonic operator, Adams’ inequality, singular exponential growth, mountain pass theorem
@@ 1. Introduction and main results
In this article, we investigate the existence of nontrivial solutions to a class of $(p,\frac{n}{2})$-biharmonic equations with singular exponential growth in $\mathbb{R}^{n}$. Specifically, we study the following equation
$\Delta^{2}_{p}u+\Delta^{2}_{\frac{n}{2}}u=\frac{g(x,u)}{|x|^{\gamma}}\quad\text{in}\quad\mathbb{R}^{n},$ ($\boldsymbol{S}$)
with $n\geq 4$, $1<p<\frac{n}{2}$, $\gamma\in(0,n)$ and $\Delta^{2}_{t}~{}\cdot:=\Delta(|\Delta\cdot|^{t-2}\Delta~{}\cdot)$ is a fourth order operator, which is known as the standard $t$-biharmonic operator for any $t\geq 1$. The nonlinearity $g:\mathbb{R}^{n}\times\mathbb{R}\to\mathbb{R}$ is a Carathéodory function and has critical exponential growth at infinity, that is, it behaves like $\exp{(\alpha|s|^{\frac{n}{n-2}})}$ as $|s|\to~{}+\infty$ for some $\alpha>0$, which means that there exists a positive constant $\alpha_{0}$ such that
uniformly with respect to $x\in\mathbb{R}^{n}$. This notion of criticality is basically driven by the well known Moser-Trudinger type inequality. Throughout this article, without further mentioning, we assume that $g:\mathbb{R}^{n}\times\mathbb{R}\to\mathbb{R}$ satisfies the following properties.
  • ($g_{1}$)
  • ($g_{2}$)
  • ($g_{3}$)
    There exists constant $\alpha_{0}>0$ such that
  • ($g_{4}$)
    There exists constant $\mu>\frac{n}{2}$ such that
  • ($g_{5}$)
    There exists positive constants $s_{0}$ and $M_{0}$ such that for all $x\in\mathbb{R}^{n}$
    $0<G(x,s)\leq M_{0}|g(x,s)|\quad\text{for all}\quad|s|\geq s_{0}.$
  • ($g_{6}$)
In addition, one can notice that, due to $(g_{2})$ and $(g_{3})$, for any fixed $q>\max\{p^{\ast},\frac{n}{2}\}$ and $\alpha>\alpha_{0}$, there exists $\zeta>0$ and a constant $D_{\zeta}>0$ depending upon $\zeta,~{}\alpha$ and $q$ such that
$|g(x,s)|\leq\zeta|s|^{\tau-1}+D_{\zeta}|s|^{q-1}\Phi_{\alpha,j_{0}}(s)\quad\text{for all}\quad(x,s)\in\mathbb{R}^{n}\times\mathbb{R}$ (1.1)
and by using $(g_{4})$, we have
$|G(x,s)|\leq\zeta|s|^{\tau}+D_{\zeta}|s|^{q}\Phi_{\alpha,j_{0}}(s)\quad\text{for all}\quad(x,s)\in\mathbb{R}^{n}\times\mathbb{R},$ (1.2)
where $\Phi_{\alpha,j_{0}}:\mathbb{R}\to[0,+\infty)$ is a Young function, which is defined by
$\Phi_{\alpha,j_{0}}(s):=\operatorname{exp}(\alpha|s|^{\frac{n}{n-2}})-\sum_{j=0}^{j_{0}-1}\frac{\alpha^{j}|s|^{\frac{nj}{n-2}}}{j!}\quad\text{with}\quad j_{0}:=\left\lceil\frac{p^{\ast}(n-2)}{n}\right\rceil=\min\Big{\{}j\in\mathbb{N}:~{}j\geq\frac{p^{\ast}(n-2)}{n}\Big{\}},$ (1.3)
[…]
Example 1.1.
We remark that the function defined by
$G(x,s):=\lambda|s|^{\vartheta}\operatorname{exp}(\alpha_{0}|s|^{\frac{n}{n-2}})\quad\text{for all}\quad(x,s)\in\mathbb{R}^{n}\times\mathbb{R}$
with $n\geq 4$, $\lambda>0$, $0<\alpha_{0}<,\alpha$, $\vartheta>\tau,\mu>\max\{p^{\ast},\frac{n}{2}\}$, $g(x,s)=\frac{\partial G(x,s)}{\partial s}$ for all $(x,s)\in\mathbb{R}^{n}\times(\mathbb{R}\setminus\{0\})$, and $g(x,0)=0$ for all $x\in\mathbb{R}^{n}$, then $g$ satisfies all the hypotheses $(g_{1})-(g_{6})$.
To deal with nonlinear elliptic PDEs, it is worthwhile to mention that several authors work with Sobolev spaces of different orders. Moreover, to understand the basic features of the Sobolev spaces, one sees that for any $m\in\mathbb{N}$ and $p\geq 1$, the space $W^{m,p}(\mathbb{R}^{n})$ can be categorized in three different ways, notably
@@
    the Gagliardo-Nirenberg-Sobolev case: $mp<n$,
@@
The Gagliardo-Nirenberg-Sobolev inequality asserts that the embedding $W^{m,p}(\mathbb{R}^{n})\hookrightarrow L^{r}(\mathbb{R}^{n})$ is continuous for all $r\in[p,\frac{np}{n-mp}]$ and $mp<n$. In this case, we can study the variational problems in the Sobolev spaces $W^{m,p}(\mathbb{R}^{n})$ with subcritical and critical polynomial growth, that is, the nonlinear term cannot exceed the polynomial of degree $\frac{np}{n-mp}$. Spite of this, the Sobolev borderline case is very special because $\frac{np}{n-mp}\to+\infty$ as $mp\to n$, which suggests us that $W^{m,p}(\mathbb{R}^{n})\subset L^{\infty}(\mathbb{R}^{n})$. Unfortunately, this only holds whenever $m=p=n=1$ and one knows it may fail for $n\geq 2$. In this instance, we note that every polynomial growth can be entertained, and thus, one may ask for a function with another kind of maximal growth. Moser-Trudinger and Adams’ inequalities play key roles in handling such difficulties. In this direction, for any bounded smooth domain $\Omega\subset\mathbb{R}^{n}~{}(n\geq 2)$ with $p=n$ and $m=1$, Trudinger [47] established that the embedding $W^{1,n}_{0}(\Omega)\hookrightarrow L_{\Phi_{\alpha}}(\Omega)$ holds, where $L_{\Phi_{\alpha}}(\Omega)$ denotes the Orlicz space associated with the Young function $\Phi_{\alpha}(t)=\operatorname{\exp}(\alpha|t|^{\frac{n}{n-1}})-1$ for some constant $\alpha>0$.The sharpness of the exponent $\alpha$ was first studied by Moser [40] and proved a version of Moser-Trudinger inequality as follows
$\sup_{u\in W^{1,n}_{0}(\Omega),~{}\|\nabla u\|_{L^{n}(\Omega)}~{}\leq 1}~{}\int_{\Omega}\operatorname{\exp}(\alpha|u|^{\frac{n}{n-1}})\,\mathrm{d}x\leq C(n)|\Omega|,~{}\forall~{}~{}\alpha\leq\alpha_{n},$ (1.4)
@@
where $\alpha_{n}=n\omega_{n-1}^{\frac{1}{n-1}}$ with $\omega_{n-1}$ is the measure of the unit sphere in $\mathbb{R}^{n}$. In addition, the above supremum is no longer finite if $\alpha>\alpha_{n}$. Later, Adimurthi-Sandeep [3] generalized the result obtained by Moser in [40] and established a sharp singular Moser-Trudinger inequality with the same Dirichlet norm used in (1.4). Both nonsingular and singular Moser-Trudinger inequalities in the Euclidean space have been studied in [12, 43] in $\mathbb{R}^{2}$, then for higher dimensions in [20, 1, 37, 4, 18, 21], respectively. Such inequalities have several applications in conformal geometry and geometric analysis of PDEs; for a detailed study, we refer to [14, 9, 30].
@@
Due to the loss of compactness in the embedding $W^{1,n}_{0}(\Omega)\hookrightarrow L_{\Phi_{\alpha}}(\Omega)$, the main difficulty arises in studying variational problems is to prove Palais-Smale compactness condition. To avoid this, Lions’ [38] introduced a concentration-compactness principle, basically a generalization of the Moser-Trudinger inequality. It states that if $u_{k}\rightharpoonup u$ weakly in $W^{1,n}_{0}(\Omega)$ with $u\neq 0$, $\|\nabla u_{k}\|_{L^{n}(\Omega)}\leq 1$, $|\nabla u_{k}|^{n}\rightharpoonup\mu$ weakly in $\mathcal{M}(\overline{\Omega})$, where $\mathcal{M}(\overline{\Omega})$ denotes the space of Radon measure in $\mathbb{R}^{n}$, then
$\sup_{k\in\mathbb{N}}\int_{\Omega}\operatorname{\exp}(\alpha_{n}\wp|u_{k}|^{\frac{n}{n-1}})\,\mathrm{d}x\leq C(\wp,\Omega),~{}\forall~{}~{}\wp\in[1,\eta),$ (1.5)
@@
where $\eta=(1-\|\nabla u^{\ast}\|^{n}_{L^{n}(\Omega)})^{-\frac{1}{n-1}}$ and $u^{\ast}$ is the spherical symmetric decreasing rearrangement of $u$. Later, in [53] the authors sharpened the constant Weierstrass-$\wp$ in (1.5) and established an improved version of Lions’ concentration-compactness principle in $W^{1,n}_{0}(\Omega)$. Similar results for $\Omega=\mathbb{R}^{n}$ can be found in [22, 50].
@@
To the best of our knowledge, there has been significant progress in the Moser-Trudinger inequalities to analyse the existence and multiplicity of solutions to quasilinear elliptic equations involving $n$-Laplace operator as well as $(p,n)$-Laplace operator in whole space $\mathbb{R}^{n}$. For a detailed study, one may see [4, 49, 29, 13].
@@
Let us introduce some notations to familiarise ourselves with higher-order Sobolev spaces. For any $m\in\mathbb{N}$ and $u\in C^{m}$, the class of $m$-th order differentiable function, we denote $\nabla^{m}$ as the $m$-th order gradient operator, which is given by
@@
Moreover, for $m<n$, the Sobolev space with homogenous Navier boundary condition on $\Omega\subset\mathbb{R}^{n}$ is denoted by $W^{m,\frac{n}{m}}_{\mathcal{N}}(\Omega)$ such that
$W^{m,\frac{n}{m}}_{\mathcal{N}}(\Omega):=\Big{\{}u\in W^{m,\frac{n}{m}}(\Omega):~{}\Delta^{j}u_{|\partial\Omega}=0~{}\text{in the sense of trace,}~{}0\leq j\leq\Big{[}\frac{m-1}{2}\Big{]}\Big{\}},$
where []delimited-[][$\cdot$][ ⋅ ] stands for greatest integer function (see Tarsi [46]). In addition, we can notice that $W^{m,\frac{n}{m}}_{0}(\Omega)\subsetneq W^{m,\frac{n}{m}}_{\mathcal{N}}(\Omega)$ and the space $W^{2,\frac{n}{2}}_{\mathcal{N}}(\Omega)$ can be explicitly defined by
$W^{2,\frac{n}{2}}_{\mathcal{N}}(\Omega):=W_{0}^{1,\frac{n}{2}}(\Omega)\cap W^{2,\frac{n}{2}}(\Omega),\quad\forall~{}n>2.$
In literature, it was Adams [2], who first studied the Moser-Trudinger inequality that appears in (1.4) into higher order Sobolev spaces. He established that for $m\in\mathbb{N}$ with $m<n$ and $\Omega$ is any bounded domain in $\mathbb{R}^{n}$, then
$\sup_{u\in W_{0}^{m,\frac{n}{m}}(\Omega)_{,}~{}\|\nabla^{m}u\|_{L^{\frac{n}{m}}(\Omega)}~{}\leq 1}~{}\int_{\Omega}\operatorname{\exp}(\beta|u|^{\frac{n}{n-m}})\,\mathrm{d}x\leq C(n,m)|\Omega|,~{}\forall~{}~{}\beta\leq\beta(n,m),$ (1.6)
@@
$\beta(n,m):=\frac{n}{\omega_{n-1}}\begin{cases}\Big{[}\frac{\pi^{\frac{n}{2}}2^{m}\Gamma\big{(}\frac{m+1}{2}\big{)}}{\Gamma\big{(}\frac{n-m+1}{2}\big{)}}\Big{]}^{\frac{n}{n-m}}\quad\text{if}~{}m~{}\text{is odd},\\ \Big{[}\frac{\pi^{\frac{n}{2}}2^{m}\Gamma\big{(}\frac{m}{2}\big{)}}{\Gamma\big{(}\frac{n-m}{2}\big{)}}\Big{]}^{\frac{n}{n-m}}\quad\quad\text{if}~{}m~{}\text{is even}.\end{cases}$
Despite this, the constant $\beta(n,m)$ is optimal in the sense that the supremum (1.6) will become infinite if $\beta>\beta(n,m)$. Moreover, (thanks to [46]) the Adams’ inequality also valid whenever $W^{m,\frac{n}{m}}(\Omega)$ is replaced by the larger space $W^{m,\frac{n}{m}}_{\mathcal{N}}(\Omega)$. Subsequently, we remark here that the above Adams’ inequality (1.6) has been extended in many directions. For example, in the bounded domain, the following inequality with singular weight has been proved in [32, Lam-Lu],
$\sup_{u\in W_{0}^{m,\frac{n}{m}}(\Omega),~{}\|\nabla^{m}u\|_{L^{\frac{n}{m}}(\Omega)}~{}\leq 1}~{}\int_{\Omega}\frac{\operatorname{\exp}(\beta|u|^{\frac{n}{n-m}})}{|x|^{\alpha}}\,\mathrm{d}x<+\infty$ (1.7)
@@
$\|u\|_{m,\frac{n}{m}}:=\begin{cases}\|(I-\Delta)^{\frac{m}{2}}u\|_{\frac{n}{m}}\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\text{if}~{}m~{}\text{is even},\\ \Big{(}\|(I-\Delta)^{\frac{m-1}{2}}u\|^{\frac{n}{m}}_{\frac{n}{m}}+\|\nabla(I-\Delta)^{\frac{m-1}{2}}u\|^{\frac{n}{m}}_{\frac{n}{m}}\Big{)}^{\frac{m}{n}}\quad\quad\quad\text{if}~{}m~{}\text{is odd},\end{cases}$
where $\|\cdot\|_{p}$ denotes $L^{p}(\mathbb{R}^{n})$ norm for $p\in[1,+\infty)$. Then the related inequalities corresponding to (1.6) when $\Omega=\mathbb{R}^{n}$ have been first proposed by [44, Ruf-Sani] for any positive even integer $m$. After that, it was again generalized in [31, Lam-Lu] for any arbitrary integer $m\geq 1$. Their combined results can be stated as follows
$\sup_{u\in W^{m,\frac{n}{m}}(\mathbb{R}^{n}),~{}\|u\|_{m,\frac{n}{m}}~{}\leq 1}~{}\int_{\mathbb{R}^{n}}\Phi(\beta(n,m)|u|^{\frac{n}{n-m}})\,\mathrm{d}x<+\infty,$ (1.8)
@@
In addition, the constant $\beta(n,m)$ in the above supremum is sharp; that is, the supremum in (1.8) is infinity if we replace $\beta(n,m)$ by any $\beta>\beta(n,m)$. Further improvement of Adams’ inequalities (1.8) have been studied in [52] and proved that (1.8) still holds under the constraint:
$\big{\{}u\in W^{m,\frac{n}{m}}(\mathbb{R}^{n}):~{}\|\nabla^{m}u\|^{\frac{n}{m}}_{\frac{n}{m}}+\tau\|u\|^{\frac{n}{m}}_{\frac{n}{m}}\leq 1~{}\text{for all}~{}\tau>0\big{\}}.$
@@
However, they did not sharpen it, and we believe it can be sharpened by utilizing a similar strategy that appears in [41, Theorem 1.1].
@@
It was Yang [48] who first extended the results in [44] and proved a version of subcritical singular Adams’ inequality in dimension four by introducing the Sobolev full norm. Basically, he established that
@@
$\sup_{u\in W^{2,2}(\mathbb{R}^{4}),~{}\int_{\mathbb{R}^{4}}(|\Delta u|^{2}+\tau|\nabla u|^{2}+\sigma|u|^{2})~{}\mathrm{d}x~{}\leq 1}~{}\int_{\mathbb{R}^{4}}\frac{\operatorname{\exp}(\alpha u^{2})-1}{|x|^{\beta}}\,\mathrm{d}x<+\infty,~{}\forall~{}\alpha\in\Big{(}0,32\pi^{2}\Big{(}1-\frac{\beta}{4}\Big{)}\Big{)}$ (1.9)
@@
for all constants $\tau,\sigma>0$ and $\beta\in[0,4)$. Further, when $\alpha>32\pi^{2}\big{(}1-\frac{\beta}{4}\big{)}$, the supremum (1.9) is no longer finite. Note that it has remained an open question to the readers whether the above inequality (1.9) still holds for the critical case, that is, $\alpha=32\pi^{2}\big{(}1-\frac{\beta}{4}\big{)}$. Then, again the singular Adams’ inequality (1.7) was fully extended into $\mathbb{R}^{n}$ if the Dirchlet norm $\|\nabla^{m}\cdot\|_{L^{\frac{n}{m}}(\Omega)}$ is replaced by $\|(\tau I-\Delta)^{\frac{m}{2}}\cdot\|_{\frac{n}{m}}$ for any $\tau>0$ (see [33]). More precisely, they proved that
@@
$\sup_{u\in W^{m,\frac{n}{m}}(\mathbb{R}^{n}),~{}\|(\tau I-\Delta)^{\frac{m}{2}}u\|_{\frac{n}{m}}~{}\leq 1}~{}\int_{\mathbb{R}^{n}}\frac{\Phi(\beta|u|^{\frac{n}{n-m}})}{|x|^{\alpha}}\,\mathrm{d}x<+\infty,~{}\forall~{}\beta\leq\beta_{\alpha,n,m},$ (1.10)

### v3

# A Detailed Analysis on Sharpened Singular Adams-Type Inequalities

@@ Abstract.
We establish a sharp Adams-type inequality in higher-order function spaces with singular weights on $\mathbb{R}^{n}.$ A sharp singular concentration–compactness principle, improving Lions’ result, is also proved. The study distinguishes between critical and subcritical sharp singular Adams-type inequalities and shows their equivalence. Furthermore, we analyze the asymptotic behavior of the associated bounds and relate the suprema of the critical and subcritical cases. A new compact embedding, crucial to our analysis, is also derived. Moreover, as an application of these results, by employing the mountain pass theorem, we study the existence of nontrivial solutions to a class of nonhomogeneous quasilinear elliptic equations involving the $(p,\frac{n}{2})$-biharmonic operator with singular exponential growth.
@@
Adams’ inequality, $(p,\frac{n}{2})$-biharmonic operator, Singular exponential growth, Mountain pass theorem
@@ 1. Introduction and main results
To deal with nonlinear elliptic PDEs, it is worth mentioning that several authors work with Sobolev spaces of different orders. Moreover, to understand the basic features of the Sobolev spaces, one sees that for any $m\in\mathbb{N}$ and $p\geq 1$, the space $W^{m,p}(\mathbb{R}^{n})$ can be categorized in three different ways, notably
@@
    the Gagliardo-Nirenberg-Sobolev (GNS, in short) case: $mp<n$,
@@
The GNS inequality asserts that the embedding $W^{m,p}(\mathbb{R}^{n})\hookrightarrow L^{r}(\mathbb{R}^{n})$ is continuous for all $r\in[p,\frac{np}{n-mp}]$ and $mp<n$. In this case, we can study the variational problems in Sobolev spaces $W^{m,p}(\mathbb{R}^{n})$ with subcritical and critical polynomial growth, that is, the nonlinear term cannot exceed the polynomial of degree $\frac{np}{n-mp}$. In spite of this, the Sobolev borderline case is very special because $\frac{np}{n-mp}\to+\infty$ as $mp\to n$, which suggests that $W^{m,p}(\mathbb{R}^{n})\subset L^{\infty}(\mathbb{R}^{n})$. Unfortunately, this only holds when $m=p=n=1$ and one knows that it may fail for $n\geq 2$. In this instance, we note that every polynomial growth can be entertained and thus, one may ask for a function with another kind of maximal growth. The Moser-Trudinger and Adams’ inequalities play key roles in handling such difficulties. In this direction, for any bounded smooth domain $\Omega\subset\mathbb{R}^{n}~(n\geq 2)$ with $p=n$ and $m=1$, Trudinger [50] established that the embedding $W^{1,n}_{0}(\Omega)\hookrightarrow L_{\Phi_{\alpha}}(\Omega)$ is valid, where $L_{\Phi_{\alpha}}(\Omega)$ denotes the Orlicz space associated with the Young function $\Phi_{\alpha}(t)=\operatorname{\exp}(\alpha|t|^{\frac{n}{n-1}})-1$ for some constant $\alpha>0$. The sharpness of the exponent $\alpha$ was first studied by Moser [43] and proved a version of the Moser-Trudinger inequality as follows
(1.1) $\sup_{u\in W^{1,n}_{0}(\Omega),~\|\nabla u\|_{L^{n}(\Omega)}~\leq 1}~\int_{\Omega}\operatorname{\exp}(\alpha|u|^{\frac{n}{n-1}})\,\mathrm{d}x\leq C(n)|\Omega|,~\forall~~\alpha\leq\alpha_{n},$
@@
where $\alpha_{n}=n\omega_{n-1}^{\frac{1}{n-1}}$ with $\omega_{n-1}$ is the measure of the unit sphere in $\mathbb{R}^{n}$. In addition, the above supremum is no longer finite if $\alpha>\alpha_{n}$. Later, Adimurthi-Sandeep [3] generalized the result obtained by Moser in [43] and established a sharp singular Moser-Trudinger inequality with the same Dirichlet norm used in (1.1). Both nonsingular and singular Moser-Trudinger inequalities in the Euclidean space have been studied in [12, 46] in $\mathbb{R}^{2}$, then for higher dimensions in [20, 1, 37, 4, 18, 21]. Such inequalities have several applications in conformal geometry and geometric analysis of PDEs; for a detailed study, we refer to [14, 9, 30].
@@
Due to the loss of compactness in the embedding $W^{1,n}_{0}(\Omega)\hookrightarrow L_{\Phi_{\alpha}}(\Omega)$, the main difficulty that arises in studying variational problems is to prove the Palais-Smale compactness condition. To avoid this, Lions’ [38] introduced a concentration-compactness principle, basically a generalization of the Moser-Trudinger inequality. It states that if $u_{k}\rightharpoonup u$ weakly in $W^{1,n}_{0}(\Omega)$ with $u\neq 0$, $\|\nabla u_{k}\|_{L^{n}(\Omega)}\leq 1$, $|\nabla u_{k}|^{n}\rightharpoonup\mu$ weakly in $\mathcal{M}(\overline{\Omega})$, where $\mathcal{M}(\overline{\Omega})$ denotes the space of the Radon measure in $\mathbb{R}^{n}$, then
(1.2) $\sup_{k\in\mathbb{N}}\int_{\Omega}\operatorname{\exp}(\alpha_{n}\wp|u_{k}|^{\frac{n}{n-1}})\,\mathrm{d}x\leq C(\wp,\Omega),~\forall~~\wp\in[1,\eta),$
@@
where $\eta=\big(1-\|\nabla u^{\ast}\|^{n}_{L^{n}(\Omega)}\big)^{-\frac{1}{n-1}}$ and $u^{\ast}$ is the spherical symmetric decreasing rearrangement of $u$. Later, in [56] the authors sharpened the constant $\wp$ in (1.2) and established an improved version of Lions’ concentration-compactness principle in $W^{1,n}_{0}(\Omega)$. Further, similar results for $\Omega=\mathbb{R}^{n}$ can be found in [22, 53].
@@
To the best of our knowledge, there has been significant progress in the Moser-Trudinger inequalities to analyzing the existence and multiplicity of solutions to quasilinear elliptic equations involving the $n$-Laplace operator as well as the $(p,n)$-Laplace operator in the whole space $\mathbb{R}^{n}$. For a detailed study, one may see [4, 52, 29, 13, 40, 41, 42] and the references therein.
@@
Let us introduce some notation to familiarize ourselves with higher-order Sobolev spaces. For any $m\in\mathbb{N}$ and $u\in C^{m}$, the class of $m$-th order differentiable function, we denote $\nabla^{m}$ as the $m$-th order gradient operator, which is given by
@@
Moreover, for $m<n$, the Sobolev space with homogeneous Navier boundary condition on $\Omega\subset\mathbb{R}^{n}$ is denoted by $W^{m,\frac{n}{m}}_{\mathcal{N}}(\Omega)$ such that
$W^{m,\frac{n}{m}}_{\mathcal{N}}(\Omega):=\Bigg\{u\in W^{m,\frac{n}{m}}(\Omega):~\Delta^{j}u_{|\partial\Omega}=0~\text{in the sense of trace,}~0\leq j\leq\Bigg[\frac{m-1}{2}\Bigg]\Bigg\},$
where $\cdot$] stands for the greatest integer function (see Tarsi [49]). In addition, we can notice that $W^{m,\frac{n}{m}}_{0}(\Omega)\subsetneq W^{m,\frac{n}{m}}_{\mathcal{N}}(\Omega)$ and the space $W^{2,\frac{n}{2}}_{\mathcal{N}}(\Omega)$ can be explicitly defined by
$W^{2,\frac{n}{2}}_{\mathcal{N}}(\Omega):=W_{0}^{1,\frac{n}{2}}(\Omega)\cap W^{2,\frac{n}{2}}(\Omega),~\forall~n>2.$
In the literature, it was Adams [2], who first studied the Moser-Trudinger inequality that appears in (1.1) into higher-order Sobolev spaces. He established that for $m\in\mathbb{N}$ with $m<n$ and $\Omega$ is any bounded domain in $\mathbb{R}^{n}$, then
(1.3) $\sup_{u\in W_{0}^{m,\frac{n}{m}}(\Omega)_{,}~\|\nabla^{m}u\|_{L^{\frac{n}{m}}(\Omega)}~\leq 1}~\int_{\Omega}\operatorname{\exp}(\beta|u|^{\frac{n}{n-m}})\,\mathrm{d}x\leq C(n,m)|\Omega|,~\forall~~\beta\leq\beta(n,m),$
@@
$\beta(n,m):=\frac{n}{\omega_{n-1}}\begin{cases}\bigg[\frac{\pi^{\frac{n}{2}}2^{m}\Gamma\big(\frac{m+1}{2}\big)}{\Gamma\big(\frac{n-m+1}{2}\big)}\bigg]^{\frac{n}{n-m}}\quad\text{if}~m~\text{is odd},\\ \bigg[\frac{\pi^{\frac{n}{2}}2^{m}\Gamma\big(\frac{m}{2}\big)}{\Gamma\big(\frac{n-m}{2}\big)}\bigg]^{\frac{n}{n-m}}\quad\quad\text{if}~m~\text{is even}.\end{cases}$
Despite this, the constant $\beta(n,m)$ is optimal in the sense that the supremum (1.3) will become infinite if $\beta>\beta(n,m)$. Moreover, (thanks to [49]) Adams’ inequality is also valid whenever $W^{m,\frac{n}{m}}(\Omega)$ is replaced by the larger space $W^{m,\frac{n}{m}}_{\mathcal{N}}(\Omega)$. Subsequently, we remark here that the above Adams’ inequality (1.3) has been extended in many directions. For example, in the bounded domain, the following inequality with singular weight has been proved in [32, Lam-Lu], which states that
(1.4) $\sup_{u\in W_{0}^{m,\frac{n}{m}}(\Omega),~\|\nabla^{m}u\|_{L^{\frac{n}{m}}(\Omega)}~\leq 1}~\int_{\Omega}\frac{\operatorname{\exp}(\beta|u|^{\frac{n}{n-m}})}{|x|^{\alpha}}\,\mathrm{d}x<+\infty$
@@
$\|u\|_{m,\frac{n}{m}}:=\begin{cases}\|(I-\Delta)^{\frac{m}{2}}u\|_{\frac{n}{m}}\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\qquad\text{if}~m~\text{is even},\\ \Big(\|(I-\Delta)^{\frac{m-1}{2}}u\|^{\frac{n}{m}}_{\frac{n}{m}}+\|\nabla(I-\Delta)^{\frac{m-1}{2}}u\|^{\frac{n}{m}}_{\frac{n}{m}}\Big)^{\frac{m}{n}}\quad\quad\qquad\text{if}~m~\text{is odd},\end{cases}$
where $\|\cdot\|_{p}$ denotes $L^{p}(\mathbb{R}^{n})$ norm for $p\in[1,+\infty)$. Then the related inequalities corresponding to (1.3) when $\Omega=\mathbb{R}^{n}$ have been first proposed by [47, Ruf-Sani] for any positive even integer $m$. After that, it was again generalized in [31, Lam-Lu] for any arbitrary integer $m\geq 1$. Their combined results can be stated as follows
(1.5) $\sup_{u\in W^{m,\frac{n}{m}}(\mathbb{R}^{n}),~\|u\|_{m,\frac{n}{m}}~\leq 1}~\int_{\mathbb{R}^{n}}\Phi(\beta(n,m)|u|^{\frac{n}{n-m}})\,\mathrm{d}x<+\infty,$
@@
In addition, the constant $\beta(n,m)$ in the above supremum is sharp; that is, the supremum in (1.5) is infinite if we replace $\beta(n,m)$ by any $\beta>\beta(n,m)$. Further improvement of Adams’ inequalities (1.5) has been studied in [55] and has shown that (1.5) still holds under the constraint
$\bigg\{u\in W^{m,\frac{n}{m}}(\mathbb{R}^{n}):~\|\nabla^{m}u\|^{\frac{n}{m}}_{\frac{n}{m}}+\tau\|u\|^{\frac{n}{m}}_{\frac{n}{m}}\leq 1~\text{for all}~\tau>0\bigg\}.$
@@
However, they did not sharpen it, and we believe it can be sharpened by using a similar strategy that used in [44, Theorem 1.1].
@@
It was Yang [51] who first extended the results in [47] and proved a version of the subcritical singular Adams’ inequality in dimension four by introducing the Sobolev full norm. Basically, he established that
@@
(1.6) $\sup_{u\in W^{2,2}(\mathbb{R}^{4}),~\int_{\mathbb{R}^{4}}(|\Delta u|^{2}+\tau|\nabla u|^{2}+\sigma|u|^{2})~\mathrm{d}x~\leq 1}~\int_{\mathbb{R}^{4}}\frac{\operatorname{\exp}(\alpha u^{2})-1}{|x|^{\beta}}\,\mathrm{d}x<+\infty,~\forall~\alpha\in\bigg(0,32\pi^{2}\bigg(1-\frac{\beta}{4}\bigg)\bigg)$
@@
for all constants $\tau,\sigma>0$ and $\beta\in[0,4)$. Furthermore, when $\alpha>32\pi^{2}\big(1-\frac{\beta}{4}\big)$, the supremum (1.6) is no longer finite. Note that it has remained an open question for the readers whether the above inequality (1.6) still holds for the critical case, that is, $\alpha=32\pi^{2}\big(1-\frac{\beta}{4}\big)$. Then again, the singular Adams’ inequality (1.4) was fully extended to $\mathbb{R}^{n}$ if the Dirichlet norm $\|\nabla^{m}\cdot\|_{L^{\frac{n}{m}}(\Omega)}$ is replaced by $\|(\tau I-\Delta)^{\frac{m}{2}}\cdot\|_{\frac{n}{m}}$ for any $\tau>0$ (see [33]). More precisely, they proved that
@@
(1.7) $\sup_{u\in W^{m,\frac{n}{m}}(\mathbb{R}^{n}),~\|(\tau I-\Delta)^{\frac{m}{2}}u\|_{\frac{n}{m}}~\leq 1}~\int_{\mathbb{R}^{n}}\frac{\Phi(\beta|u|^{\frac{n}{n-m}})}{|x|^{\alpha}}\,\mathrm{d}x<+\infty,~\forall~\beta\leq\beta_{\alpha,n,m},$